OpenStudy (andromeda992):

Please Help! A stone falls on a ledge and takes 8 seconds to hit the ground. The stone has an original velocity of 0 m/s. How tall is the ledge? A. 39.2 meters B. 78.4 meters C. 1254.4 meters D 13.58 meters

OpenStudy (agent0smith):

http://www.physicsclassroom.com/Class/1DKin/U1L6a1.gif Top right equation

OpenStudy (mrnood):

s= ut + (at^2)/2 u=0 t= 8 a=9.8 work out s for th eheight

OpenStudy (mrnood):

have you written this correctly? - none of th eanswers seem correct to me

OpenStudy (andromeda992):

yes i did

OpenStudy (mrnood):

@agent0smith it is not tgh etop right equation - they do not want the final velocity - they weant th edistance top left equation. But none of the given answers fits the question

OpenStudy (agent0smith):

Top left.

OpenStudy (agent0smith):

You knew what i meant!

OpenStudy (mrnood):

I knew - but th eposter didn't

OpenStudy (agent0smith):

Well, I knew what I meant.

OpenStudy (andromeda992):

Okay so then I don't use the equation @MrNood posted

OpenStudy (agent0smith):

I didn't even notice, lol And no answers are correct

OpenStudy (agent0smith):

The building should be approx. 320 m high.

OpenStudy (andromeda992):

Okay so I just realized I did put down the wrong thing

OpenStudy (mrnood):

using d= ut + 0.5 a t^2 u=0 a=9.81 t=8 d=0 + 9.81 *64/2 =314 m none of the answers are correct - check again thast you wrote the question correctlyy

OpenStudy (andromeda992):

D. is actually 313.6 m

OpenStudy (andromeda992):

*lowers head in shame* Sorry

OpenStudy (mrnood):

forgiven :-)

OpenStudy (mrnood):

BTW - Empire State Building is 380m

OpenStudy (andromeda992):

lol thanks